Part 1, 2, & 3
Check my work During May. Joliet Fabrics Corporation manuf nformation from the Stylex production department also pertains t octured 620 unts floyer fabric with the trade name Stylex The tollowing Direet material purchased: 19,200 yards at $1.50 per yard Direct material used: 10,700 yards at $1.50 per yard Direct labor: 3,300 hours at $9.27 per hour $ 28,800 16,050 30,591 The standard prime costs for one unit of Stylex are as follows Direct material: 20 yards at $1.47 per yard Direct labor: 4 hours at $9.00 per hour $29.40 36.00 Total standard prime cost per unit of output-$65.10 Required: Compute the following variances for the month of May. (Indicate the effect of each variance by selecting Favorable or Unfavorable. Select None and enter O for no effect (i.e., zero variance). 1. Direct-material price variance Unfavorable
media%2F731%2F7313f1fb-a6b1-420c-83c9-68
media%2Fca4%2Fca4af294-6dbd-43bc-9d74-4c
media%2Fccb%2Fccbd5c9a-047b-42c5-a949-fd
media%2F879%2F879ac82f-5215-454d-95d6-c2
Answer #1

Solution 1:

Standard quantity of material for actual production = 620 * 20 = 12400 yards

Actual quantity of material purchased = 19200 yards

Actual quantity of material consumed = 10700 yards

Standard price of material = $1.47 per yard

Actual price of material = $1.50 per yard

Material price variance = (SP - AP) * AQ used = ($1.47 - $1.50) * 10700 = $321 U

Material quantity variance = (SQ - AQ) * SP = (12400 - 10700) * $1.47 = $2,499 F

Material purchase price variance = (SP - AP) * AQ Purchased = ($1.47 - $1.50) * 19200 = $576 U

Standard hours of direct labor for actual production = 620 * 4= 2480 hours

Actual hours of direct labor = 3300 hours

Standard rate of direct labor = $9 per hour

Actual rate of direct labor = $9.27 per hour

Direct labor rate variance = (SR - AR) * AH = ($9 - $9.27) * 3300 = $891 U

Direct labor efficiency variance = (SH - AH) * SR = (2480 - 3300) * $9 = $7,380 U